OpenStudy (anonymous):

which equation is related equation for this equation 23/a=8 A. A+8=23 B. A=23x8 C. A=23/8 D. A=8/23

OpenStudy (anonymous):

c

OpenStudy (anonymous):

help me with one more plz

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Which value is a solution to the open sentence? 6x – 3 = 39 A. 4 B. 6 C. 5 D. 7

OpenStudy (anonymous):

sorry wrong one

OpenStudy (anonymous):

ok which one do u need help with

OpenStudy (anonymous):

What is the solution to the equation? 5x = 80 A. x = 16 B. x = 75 C. x = 85 D. x = 400

OpenStudy (anonymous):

to solve to slove this you take the question 5x =80 and you divide both sides by 5x so 5x=80=5x/5x=x then 80/5=16x=16

OpenStudy (anonymous):

x=16

OpenStudy (anonymous):

so it's A?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What is the solution to the equation? 5x = 80 A. x = 16 B. x = 75 C. x = 85 D. x = 400

OpenStudy (anonymous):

a) x=16 is the answer

OpenStudy (anonymous):

Which word phrase represents the variable expression? 2n + 4 A. twice a number decreased by 4 B. four more than twice a number C. two more than 4 times a number D. twice the sum of a number and 4